OpenStudy (anonymous):

if y=x^3/z^2,what happens to the value of y when both x and z are doubled? A)y does not change B)y is halved c)y is doubled d)y is multiplied by 4 e)y is multiplied by 8

OpenStudy (anonymous):

E i am sure Here is how you go about solving it: y=5(2x)^3/2z, i.e, the values of x and z are multiplied by 2, Then: y=5*8*x^3/2z y=(5x^3/z)*(8/2)=(5x^3/z)*4

OpenStudy (asnaseer):

replace x with 2x and z with 2z in the expression and see what you end up with